sec of theta= -6 given that sin theta is > 0. find the remaining six trigonometric functions of theta
Answer by stanbon(75887)

sec of theta= -6 given that sin theta is > 0. find the remaining six trigonometric functions of theta
----------
Since sec is negative and sin is positive theta is in QII
By definition sec = r/x
Since sec(t) = -6, r = 6 and x = -1
Then y = sqrt[r^2-x^2[ = sqrt(36-1] = sqrt(35)
=====
Ans:: sin(t) = y/r = sqrt(35)/6
-------------------
Cheers,
Stan H.